If three times the smaller of two consecutive integers is added to four times the larger, the result is 88. Find the smaller integer.
Answer by stanbon(75887)

If three times the smaller of two consecutive integers is added to four times the larger, the result is 88. Find the smaller integer.
----------------------------
Let smaller be x
Then larger is x+1
--------------------------------------
Equation:
3x+4(x+1) = 88
3x+4x+4 = 88
7x = 84
x = 12 (the smaller)
x+1 = 13 (the larger)
==========================
Cheers,
Stan H.
======================